# The Potential of Second-Order Optimization for LLMs: A Study with Full Gauss-Newton

### v1

@@ 1 Introduction
With rising compute requirements for training large language models (LLMs), improving optimization methods has become a central strategy for improving training efficiency. Better optimizers can directly reduce the serial runtime to train an LLM, which is crucial for large-scale models that train from days to months. Optimization for LLMs has traditionally leveraged first-order methods such as SGD and Adam (Kingma & Ba, 2017). However, recent research in optimization has started exploring the use of second-order optimizers for large-scale models, motivated by the faster convergence rates known from theory (Nesterov, 2018) and potential to scale to larger batch sizes (Zhang et al., 2019) – two ways of reducing serial runtime.
@@
To answer this, we first establish the performance limits of an idealized second-order method, full Gauss-Newton (GN), and measure its performance in terms of iteration complexity (the number of steps to reach a target loss). This serves as a practical lower bound for any second-order approach. We also analyze how this idealized method affects the critical batch size (McCandlish et al., 2018; Shallue et al., 2019; Jain et al., 2018b), a key measure of data parallelism efficiency.
@@ 2 Related work
We mention a few of the most related works here and provide additional related work in Appendix B; work on specific optimizers for LLMs is discussed in Section 3. Most related to our work is Hessian-free optimization, which avoids explicit Hessian formation by leveraging Hessian-vector products (Martens, 2010). This approach serves as an alternative to layerwise approximation methods of the Hessian as discussed in Section 3. Specifically, prior work on Hessian-free optimizers use the conjugate gradient (CG) to solve an incomplete (unconverged) optimization of the Newton step rather than storing an approximation to the Hessian. This is introduced by Martens (2010) on classification and auto-encoder tasks, and extended to additional settings such as recurrent neural networks by Martens & Sutskever (2011a) and Cho et al. (2015). Garcia et al. (2023) amortizes the CG steps in Hessian-free optimization for deep linear and auto-encoder models. In contrast, our work focuses on the setting of LLMs, and we leverage optimizers that are specifically designed for LLMs (e.g. Adam and Muon) rather than CG to apply the Gauss-Newton step.
@@ 3 Background on existing optimizers
The most widely used optimizer for LLMs is Adam (Kingma & Ba, 2017). Adam maintains matrices for the first and second moment of the gradient $G_{t}$, denoted $M_{t}$ and $V_{t}$ respectively. Adam performs the element-wise update
@@
SOAP (Vyas et al., 2025) is a recent variant of Shampoo that runs AdamW (Loshchilov & Hutter, 2019) in the eigenbasis provided by Shampoo.
@@
where $NS$ denotes a Newton-Schulz orthogonalization procedure. (See Jordan et al. (2024b) for further description of the Newton-Schulz method). Muon without momentum can be seen as a version of Shampoo without preconditioner accumulation (Bernstein & Newhouse, 2024).
@@ 4.2 Newton’s method & the Gauss-Newton matrix
That is, the Gauss-Newton matrix is defined as $G\coloneqq\nabla_{\theta}f(\theta)^{\mathsf{T}}\nabla_{z}^{2}\mathcal{L}(\theta)\nabla_{\theta}f(\theta)$. Intuitively, the Gauss-Newton captures the curvature of the loss function, but drops the curvature of the model. Unlike the Hessian, the Gauss-Newton matrix is PSD for MSE and cross-entropy loss (Martens, 2020). This avoids untrustworthy updates as negative curvature implies unbounded decrease in loss (Martens, 2020). Indeed, methods using the Gauss-Newton matrix rather than the full Hessian have been found to lead to better optimization (Martens, 2010; Martens & Sutskever, 2012; Vinyals & Povey, 2012).
@@ 4.3 Memory-feasible Gauss-Newton implementation
However, given that computing the Gauss-Newton matrix directly is infeasible, we instead run a functionally equivalent method that leverages Jacobian-vector products (JVPs) to avoid explicitly storing the Hessian. Specifically, we optimize the second-order Taylor approximation of the loss function $\mathcal{L}$ with a first-order Taylor approximation of the model $f$. The minimization of the loss in this setting is equivalent to using the Gauss-Newton matrix as a preconditioner (Martens & Sutskever, 2011b). The proof is provided in Appendix A.
@@ 5 Experiment details
We train 45M and 150M parameter LLaMA models (Touvron et al., 2023) on the C4 dataset (Raffel et al., 2020). Full details on models and hyperparameter sweeps are given in Appendix D and Appendix G respectively.
@@
We run AdamW (Loshchilov & Hutter, 2019), Muon (Jordan et al., 2024b), and SOAP (Vyas et al., 2025) as baselines. For 45M models, for each method at each batch size, we run a hyperparameter sweep over learning rate, weight decay, and weight averaging decay if applicable. We additionally sweep the $\beta_{2}$ parameter for Adam, the $\mu$ parameter for Muon, and $(\beta_{1},\beta_{2})$ for SOAP. For 150M parameter models we run a more limited hyperparameter sweep over learning rate and $\beta$ and $\mu$ parameters. To make sure runs are well-initialized, we start all runs after an AdamW warmup consisting of $5\%$ of the Chinchilla-optimal number of tokens (Hoffmann et al., 2022). More details on hyperparameter tuning are given in Appendix G.
@@
For each training step, we take a first-order Taylor approximation of the model around the current parameters. We initialize the parameters of the Taylor approximation to be the pre-linesearch parameters from the previous iteration (see Section 5.1). We then take a second-order Taylor approximation around the cross-entropy loss on the Taylorized model, also around the current model parameters. We use Muon (Jordan et al., 2024b) with batch size $b_{inner}$ to minimize the Taylorized loss. We take $N$ steps of Muon and then update the model parameters using a line search. We refer to the global batch size (in number of sequences) as $b=N\times b_{inner}$ since this is the total amount of data seen per parameter update on the true model. We use $b_{inner}=32$ for the 45M models and $b_{inner}=128$ for the 150M models with sequence length 1024, and vary $N$ to control the overall batch size $b$. We start all runs from the same AdamW post-warmup checkpoint. To compute the necessary Taylor approximations we use the neural-tangents library from Novak et al. (2020). See Algorithm 1 for details.
@@ 6.1.1 Iteration complexity
We find that the Gauss-Newton method can make fast progress in the large batch size regime, particularly in the first few steps of training. After 10 steps, the loss for the Gauss-Newton model is below 3.75, while other methods have made marginal progress from the starting loss. The optimal learning rates for AdamW, Muon, and SOAP lead to initial instability but faster convergence overall – see Appendix G for details on choosing hyperparameters. The Gauss-Newton method is able to reach loss 3.25 in 54 steps, a 5.4x gain over SOAP and 16x gain over Muon. The results are shown in Figure 1.
@@ 6.1.2 Batch size scaling
While iteration complexity captures a purely sequential perspective of training efficiency, it is also important to consider the sample efficiency. In an ideal setting, the number of samples seen at each step would vary proportionally to the number of steps, such that there is always a constant number of samples required overall. However, it is known that sample efficiency is lost once the batch size is scaled past a given method’s critical batch size (McCandlish et al., 2018; Shallue et al., 2019). That is, the total number of samples needed to achieve a given loss will grow once the critical batch size is exceeded. Therefore, we study the batch size scaling behavior of the Gauss-Newton method to understand how much we lose in sample and computational efficiency when we minimize the number of sequential iterations. We run experiments in two settings: first, we train models for a fixed amount of data over a range of batch sizes and measure the final loss. Second, we measure how the number of steps to achieve a given loss changes with increasing batch size. We run experiments for 45M- and 150M-parameter models; see Appendix F for results at 45M-parameter scale.
@@
Following the methodology of Zhang et al. (2025), we plot the number of steps required for each optimization method to reach the target validation loss of 3.4 as a function of batch size. The point at which the curve for each model plateaus defines its respective critical batch size (McCandlish et al., 2018; Shallue et al., 2019; Jain et al., 2018b). We find that AdamW levels off near batch size of 4M with little further reduction. SOAP and Muon continue to decrease up to batch size of 12M but with diminishing reductions, and show little additional decrease by 40M. Meanwhile, the Gauss-Newton method continues to decrease through 40M, indicating better sample efficiency at large batch sizes.
@@ 6.2 GN-prox-linear method
See Algorithm 2 for details. Note that for any convex loss function (including cross-entropy loss), the above optimization problem is still convex. Moreover, this problem is related to the richly studied literature of kernelized classification (Shalev-Shwartz & Ben-David, 2014) (albeit with cross-entropy loss, instead of the max-margin loss). This GN-prox-linear method is notably not a second-order method. Rather, it allows us to study the effect of the higher order terms of the loss as compared to the Gauss-Newton update rule.
We train in the same Chinchilla-optimal setting on 150M parameter models and perform the same hyperparameter sweeps as for Gauss-Newton (See Appendix G). We find that the inclusion of higher order loss terms has little effect on performance as compared to Gauss-Newton; results are shown in Figure 3. However, unlike Gauss-Newton, we found that the global cosine schedule for the inner optimizer outperformed the constant+inner cosine schedule. Additionally, line search was not necessary for the GN-prox-linear method.
@@ 6.3 Layerwise Gauss-Newton
Due to compute costs, we train 150M parameter layerwise Gauss-Newton models only for the largest three batch sizes. We follow the same training setting for fixed token count as specified in Section 6.1.2. For layerwise experiments we set hyperparameters to match those of the best Gauss-Newton configuration at each batch size. However, we include smaller step size options for the line search as we find this is necessary for stable convergence. We provide more details on line search step sizes in Appendix C. We find that the layerwise Gauss-Newton method also achieves comparable performance through batch size of 40M tokens (Figure 3). We additionally train a layerwise Gauss-Newton model with batch size of 120M tokens to loss 3.25 to compare its iteration complexity to that of full Gauss-Newton (See Sec 6.1.1). We find that the layerwise Gauss-Newton takes only 1.4x more steps to reach the target loss compared to the full Gauss-Newton method and provides a 3.4x gain over SOAP (Figure 1).
@@ 7 Discussion and Conclusion
In this work, we study whether full second-order optimization – specifically, using the full Gauss-Newton matrix as a preconditioner – can offer further benefits for training large language models as compared to existing methods. In particular, we focus on the large batch size regime, following Jain et al. (2018a) and Zhang et al. (2019) which show that the benefits of preconditioning may not appear at small batch size. While our current implementation is roughly 4-5x slower than standard training (e.g. with AdamW or Muon), we view this as a proof of concept demonstrating the potential of exact second-order methods. Our results indicate that further development in second-order methods could lead to substantial benefits in convergence and ability to scale to larger batch size.
@@ Appendix A Proof of equivalence to Gauss-Newton method
Since the Gauss-Newton matrix is PSD, we can find $\theta^{*}$ to minimize $h^{(2)}$ by setting its gradient to zero:
@@
Appendix B Additional related work
@@
Appendix C Additional details on optimization strategies
C.1 Learning rate schedules
@@
C.2 Line search step sizes
@@
Appendix D Model details
@@
Appendix E Compute resources
@@
Appendix F Experiments on 45M parameter models
@@
Appendix G Details on hyperparameter tuning

### v2

@@ Abstract
Code is available at: https://github.com/natalieabreu/full-gauss-newton.
@@ 1 Introduction
With rising compute requirements for training large language models (LLMs), improving optimization methods has become a central strategy for improving training efficiency. Better optimizers can directly reduce the serial runtime to train an LLM, which is crucial for large-scale models that train from days to months. Optimization for LLMs has traditionally leveraged first-order methods such as SGD and Adam (Kingma and Ba, 2017). However, recent research in optimization has started exploring the use of second-order optimizers for large-scale models, motivated by the faster convergence rates known from theory (Nesterov, 2018) and potential to scale to larger batch sizes (Zhang et al., 2019) – two ways of reducing serial runtime.
@@
To answer this, we first establish the performance limits of an idealized second-order method, full Gauss-Newton (GN), and measure its performance in terms of iteration complexity (the number of steps to reach a target loss). This serves as a practical lower bound for any second-order approach. We also analyze how this idealized method affects the critical batch size (McCandlish et al., 2018; Shallue et al., 2019; Jain et al., 2018a), a key measure of data parallelism efficiency.
@@ 2 Related work
We mention a few of the most related works here and provide additional related work in Appendix C; work on specific optimizers for LLMs is discussed in Section 3. Most related to our work is Hessian-free optimization, which avoids explicit Hessian formation by leveraging Hessian-vector products (Martens, 2010). This approach serves as an alternative to layerwise approximation methods of the Hessian as discussed in Section 3. Specifically, prior work on Hessian-free optimizers use the conjugate gradient (CG) to solve an incomplete (unconverged) optimization of the Newton step rather than storing an approximation to the Hessian. This is introduced by Martens (2010) on classification and auto-encoder tasks, and extended to additional settings such as recurrent neural networks by Martens and Sutskever (2011a) and Cho et al. (2015). Garcia et al. (2023) amortizes the CG steps in Hessian-free optimization for deep linear and auto-encoder models. In contrast, our work focuses on the setting of LLMs, and we leverage optimizers that are specifically designed for LLMs (e.g. Adam and Muon) rather than CG to apply the Gauss-Newton step.
@@ 3 Background on existing optimizers
The most widely used optimizer for LLMs is Adam (Kingma and Ba, 2017). Adam maintains matrices for the first and second moment of the gradient $G_{t}$, denoted $M_{t}$ and $V_{t}$ respectively. Adam performs the element-wise update
@@
SOAP (Vyas et al., 2025) is a recent variant of Shampoo that runs AdamW (Loshchilov and Hutter, 2019) in the eigenbasis provided by Shampoo.
@@
where $NS$ denotes a Newton-Schulz orthogonalization procedure. (See Jordan et al. (2024b) for further description of the Newton-Schulz method). Muon without momentum can be seen as a version of Shampoo without preconditioner accumulation (Bernstein and Newhouse, 2024).
@@ 4.2 Newton’s method & the Gauss-Newton matrix
That is, the Gauss-Newton matrix is defined as $G\coloneqq\nabla_{\theta}f(\theta)^{\mathsf{T}}\nabla_{z}^{2}\mathcal{L}(\theta)\nabla_{\theta}f(\theta)$. Intuitively, the Gauss-Newton captures the curvature of the loss function, but drops the curvature of the model. Unlike the Hessian, the Gauss-Newton matrix is PSD for MSE and cross-entropy loss (Martens, 2020). This avoids untrustworthy updates as negative curvature implies unbounded decrease in loss (Martens, 2020). Indeed, methods using the Gauss-Newton matrix rather than the full Hessian have been found to lead to better optimization (Martens, 2010; Martens and Sutskever, 2012; Vinyals and Povey, 2012).
@@ 4.3 Memory-feasible Gauss-Newton implementation
However, given that computing the Gauss-Newton matrix directly is infeasible, we instead run a functionally equivalent method that leverages Jacobian-vector products (JVPs) to avoid explicitly storing the Hessian. Specifically, we optimize the second-order Taylor approximation of the loss function $\mathcal{L}$ with a first-order Taylor approximation of the model $f$. The minimization of the loss in this setting is equivalent to using the Gauss-Newton matrix as a preconditioner (Martens and Sutskever, 2011b). The proof is provided in Appendix A.
@@ 5 Experiment details
We train 45M and 150M parameter LLaMA models (Touvron et al., 2023) on the C4 dataset (Raffel et al., 2020). Full details on models and hyperparameter sweeps are given in Appendix E and Appendix H respectively.
@@
We run AdamW (Loshchilov and Hutter, 2019), Muon (Jordan et al., 2024b), and SOAP (Vyas et al., 2025) as baselines. For 45M models, for each method at each batch size, we run a hyperparameter sweep over learning rate, weight decay, and weight averaging decay if applicable. We additionally sweep the $\beta_{2}$ parameter for Adam, the $\mu$ parameter for Muon, and $(\beta_{1},\beta_{2})$ for SOAP. For 150M parameter models we run a more limited hyperparameter sweep over learning rate and $\beta$ and $\mu$ parameters. To make sure runs are well-initialized, we start all runs after an AdamW warmup consisting of $5\%$ of the Chinchilla-optimal number of tokens (Hoffmann et al., 2022). More details on hyperparameter tuning are given in Appendix H.
@@
For each training step, we take a first-order Taylor approximation of the model around the current parameters. We initialize the parameters of the Taylor approximation to be the pre-linesearch parameters from the previous iteration (see Section 5.1). We then take a second-order Taylor approximation around the cross-entropy loss on the Taylorized model, also around the current model parameters. We use Muon (Jordan et al., 2024b) with batch size $b_{inner}$ to minimize the Taylorized loss. We take $N$ steps of Muon and then update the model parameters using a line search. We refer to the global batch size (in number of sequences) as $b=N\times b_{inner}$ since this is the total amount of data seen per parameter update on the true model. We use $b_{inner}=32$ for the 45M models and $b_{inner}=128$ for the 150M models with sequence length 1024, and vary $N$ to control the overall batch size $b$. We start all runs from the same AdamW post-warmup checkpoint. To compute the necessary Taylor approximations we use the neural-tangents library from Novak et al. (2020). See Algorithm 1 for details. In addition we provide an efficient implementation for the minimization of the least squares subproblem in Appendix B.
@@ 6.1.1 Iteration complexity
We find that the Gauss-Newton method can make fast progress in the large batch size regime, particularly in the first few steps of training. After 10 steps, the loss for the Gauss-Newton model is below 3.75, while other methods have made marginal progress from the starting loss. The optimal learning rates for AdamW, Muon, and SOAP lead to initial instability but faster convergence overall – see Appendix H for details on choosing hyperparameters. The Gauss-Newton method is able to reach loss 3.25 in 54 steps, a 5.4x gain over SOAP and 16x gain over Muon. The results are shown in Figure 1.
@@ 6.1.2 Batch size scaling
While iteration complexity captures a purely sequential perspective of training efficiency, it is also important to consider the sample efficiency. In an ideal setting, the number of samples seen at each step would vary proportionally to the number of steps, such that there is always a constant number of samples required overall. However, it is known that sample efficiency is lost once the batch size is scaled past a given method’s critical batch size (McCandlish et al., 2018; Shallue et al., 2019). That is, the total number of samples needed to achieve a given loss will grow once the critical batch size is exceeded. Therefore, we study the batch size scaling behavior of the Gauss-Newton method to understand how much we lose in sample and computational efficiency when we minimize the number of sequential iterations. We run experiments in two settings: first, we train models for a fixed amount of data over a range of batch sizes and measure the final loss. Second, we measure how the number of steps to achieve a given loss changes with increasing batch size. We run experiments for 45M- and 150M-parameter models; see Appendix G for results at 45M-parameter scale.
@@
Following the methodology of Zhang et al. (2025), we plot the number of steps required for each optimization method to reach the target validation loss of 3.4 as a function of batch size. The point at which the curve for each model plateaus defines its respective critical batch size (McCandlish et al., 2018; Shallue et al., 2019; Jain et al., 2018a). We find that AdamW levels off near batch size of 4M with little further reduction. SOAP and Muon continue to decrease up to batch size of 12M but with diminishing reductions, and show little additional decrease by 40M. Meanwhile, the Gauss-Newton method continues to decrease through 40M, indicating better sample efficiency at large batch sizes.
@@ 6.2 GN-prox-linear method
See Algorithm 2 for details. Note that for any convex loss function (including cross-entropy loss), the above optimization problem is still convex. Moreover, this problem is related to the richly studied literature of kernelized classification (Shalev-Shwartz and Ben-David, 2014) (albeit with cross-entropy loss, instead of the max-margin loss). This GN-prox-linear method is notably not a second-order method. Rather, it allows us to study the effect of the higher order terms of the loss as compared to the Gauss-Newton update rule.
We train in the same Chinchilla-optimal setting on 150M parameter models and perform the same hyperparameter sweeps as for Gauss-Newton (See Appendix H). We find that the inclusion of higher order loss terms has little effect on performance as compared to Gauss-Newton; results are shown in Figure 3. However, unlike Gauss-Newton, we found that the global cosine schedule for the inner optimizer outperformed the constant+inner cosine schedule. Additionally, line search was not necessary for the GN-prox-linear method.
@@ 6.3 Layerwise Gauss-Newton
Due to compute costs, we train 150M parameter layerwise Gauss-Newton models only for the largest three batch sizes. We follow the same training setting for fixed token count as specified in Section 6.1.2. For layerwise experiments we set hyperparameters to match those of the best Gauss-Newton configuration at each batch size. However, we include smaller step size options for the line search as we find this is necessary for stable convergence. We provide more details on line search step sizes in Appendix D. We find that the layerwise Gauss-Newton method also achieves comparable performance through batch size of 40M tokens (Figure 3). We additionally train a layerwise Gauss-Newton model with batch size of 120M tokens to loss 3.25 to compare its iteration complexity to that of full Gauss-Newton (See Sec 6.1.1). We find that the layerwise Gauss-Newton takes only 1.4x more steps to reach the target loss compared to the full Gauss-Newton method and provides a 3.4x gain over SOAP (Figure 1).
@@ 7 Discussion and Conclusion
In this work, we study whether full second-order optimization – specifically, using the full Gauss-Newton matrix as a preconditioner – can offer further benefits for training large language models as compared to existing methods. In particular, we focus on the large batch size regime, following Jain et al. (2018b) and Zhang et al. (2019) which show that the benefits of preconditioning may not appear at small batch size. While our current implementation is roughly 1.5x slower than standard training (e.g. with AdamW or Muon), we view this as a proof of concept demonstrating the potential of exact second-order methods. Our results indicate that further development in second-order methods could lead to substantial benefits in convergence and ability to scale to larger batch size.
@@ Appendix A Proof of equivalence to Gauss-Newton method
[…]
@@
Appendix B Implementation of Gauss-Newton
Following notation from Appendix A, we have the Gauss-Newton objective:
[…]
Then the gradient is:
[…]
We compute the gradient directly in the following manner, which results in only one additional forward pass as compared to a standard backpropagation step:
Input: train_state $(\theta)$,
reference params $\theta_{0}$,
batch $\mathcal{B}=\{\text{input},\text{target}\}$
Define $f_{\text{batch}}(\theta)$:
$\text{logits}=f(\theta,\;\mathcal{B.\text{input}})$
return logits
// $f\text{ on current batch }\mathcal{B}$
Define $\ell_{\text{logits}}(y)$:
$\text{loss}\leftarrow\text{CE\_loss}\big(y,\;\mathcal{B}.\text{target})$
return loss
Define $\textsc{ValueAndGradient}(\theta_{0},\theta)$:
// Linearize $f$ at $\theta_{0}$: $f(\theta)\approx y_{0}+J_{0}(\theta-\theta_{0})$
$(y_{0},\;\mathrm{jvp})\leftarrow\text{linearize}(f_{\text{batch}},\;\theta_{0})$
$d\theta\leftarrow\theta-\theta_{0}$
$v\leftarrow\mathrm{jvp}(d\theta)$
// $v=J_{0}\,d\theta$
1ex
// Logit-space gradient and Hessian-vector
$\nabla_{y}\ell_{\text{logits}}\leftarrow\text{grad}(\ell_{\text{logits}})\;$
$(g_{0},\;H_{y}v)\leftarrow\text{JVP}\big(\nabla_{y}\ell_{\text{logits}},\;(y_{0}),\;(v)\big)$
// $H_{y}=\nabla^{2}_{yy}\ell$ at $y_{0}$
1ex
// Pullback through the linearization: $J_{0}^{\top}(g_{0}+H_{y}v)$
$\mathrm{jt}\leftarrow\text{linear\_transpose}(\mathrm{jvp},\;\theta_{0})$
$\nabla_{\theta}\leftarrow\mathrm{jt}(g_{0}+H_{y}v)$
1ex
// Quadratic objective on the linear model
$\text{loss}\leftarrow\ell_{\text{logits}}(y_{0})\;+\;g_{0}\cdot v\;+\;\frac{1}{2}\,v\cdot(H_{y}v)$
return $\big(\text{loss},\;\nabla_{\theta}\big)$
$\text{loss},\;g\leftarrow\textsc{ValueAndGradient}(\theta_{0},\;\theta=\text{train\_state.params})$
1ex
Apply update:
$\text{train\_state}\leftarrow\text{train\_state.apply\_gradients}(g)$
1ex
return train_state
Algorithm 3 Gauss-Newton Inner Optimization Step
Appendix C Additional related work
@@
Appendix D Additional details on optimization strategies
D.1 Learning rate schedules
@@
D.2 Line search step sizes
@@
Appendix E Model details
@@
Appendix F Compute resources
@@
Appendix G Experiments on 45M parameter models
@@
Appendix H Details on hyperparameter tuning